OpenStudy (anonymous):

Which expression has a value of –12? A. –4 – (–8) B. 3 – 9 C. –16 – (–4) D. –6 – 18

OpenStudy (alexandervonhumboldt2):

-4-(-8)=-4+8=4 3-9=-6 -16-(-4)=-16+4=-12 -6-18=-24

OpenStudy (alexandervonhumboldt2):

-(-)=+

OpenStudy (anonymous):

so it is C

OpenStudy (alexandervonhumboldt2):

yeah